OpenStudy (anonymous):

If n > 1 and y = log n x, find dy/dx.

OpenStudy (anonymous):

Did you mean \(y=\log(nx)\)?

OpenStudy (anonymous):

If so, recall that the derivative of \(\ln(x)\) is \(\frac{1}{x}\) and use the chain rule.

OpenStudy (anonymous):

(Assuming you're using \(\log\) to mean the natural logarithm, not base 10.)

OpenStudy (anonymous):

ah, n is actually the base! \[\log_{n}x\]

OpenStudy (anonymous):

Oh, then the rule to use is \(\log_a(b)^{\prime} = \frac{1}{x\ln(a)}\).

OpenStudy (anonymous):

I started with a b and ended with an x there, sorry.

OpenStudy (anonymous):

So in your case you would get \[\frac{1}{x\ln(n)}\]

OpenStudy (anonymous):

okay, i understand now! thank you. the n instead of a constant confused me.

OpenStudy (anonymous):

You can treat any variables besides the one you're taking a derivative with respect to as constants.

OpenStudy (anonymous):

(As long as they aren't functions of your variable, \(y(x)\) obviously isn't a constant.)